OpenStudy (anonymous):

f(x)=1/5x+1, then find f(20). any know fractions?

OpenStudy (anonymous):

\[f(x) = \frac{1}{5x} + 1\] or \[f(x) = \frac{1}{5}x+ 1\] which one?

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

y=1/5(20)+1 y=4+1 y=5 OR y=1/5(20) + 1 y=1/100+1 y=1.01

OpenStudy (anonymous):

If it's the second one, then y = 5

OpenStudy (anonymous):

im still lost! what do u do with the fraction?

OpenStudy (anonymous):

\[\frac{1}{5}* 20 + 1 = \frac{20}{5} + 1\] just divide 20 by 5 to get \[4 + 1 = 5\]

OpenStudy (anonymous):

ok